OpenStudy (anonymous):

what is the slope of a line with the equation y=5x+2/5

OpenStudy (anonymous):

It is of the form y=mx+c where m is the slope and c is the y-intercept. Hence, slope is 5.

OpenStudy (anonymous):

If you have an equation in the form of y = mx + b The slope is the m. The number in front of the x.

OpenStudy (anonymous):

the slope is 5 because its already in y=mx+b form and i the slope is m. so its 5.

OpenStudy (anonymous):

5 as simple as it can get